OpenStudy (didierr):

Which of the following is equal to 7 1/2 ?

OpenStudy (alexandervonhumboldt2):

\[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\]

OpenStudy (alexandervonhumboldt2):

\[7\frac{ 1 }{ 2 }=\frac{ 7*2+1 }{ 2}=\]

OpenStudy (didierr):

A 13/2 B 4 9/4 C 5 5/2 D 25/4

OpenStudy (alexandervonhumboldt2):

i showed you in main form how to convert the fractions, can you please now convert it yourself?

OpenStudy (didierr):

yea sure